# An Invitation to Deep Reinforcement Learning

### v1

@@ 1 Introduction
The field of reinforcement learning (RL) is traditionally viewed as the art of learning by trial and error (Sutton & Barto, 2018). Reinforcement learning methods were historically developed to solve sequential decision making tasks. The core idea is to deploy an untrained model in an environment. This model is called the policy and maps inputs to actions. The policy is then improved by randomly attempting different actions and observing an associated feedback signal, called the reward. Reinforcement learning techniques have demonstrated remarkable success when applied to popular games. For example, RL produced world-class policies in the games of Go (Silver et al., 2016; 2018; Schrittwieser et al., 2020), Chess (Silver et al., 2018; Schrittwieser et al., 2020), Shogi (Silver et al., 2018; Schrittwieser et al., 2020), Starcraft (Vinyals et al., 2019), Stratego (Perolat et al., 2022) and achieved above human level policies in all Atari games (Badia et al., 2020; Ecoffet et al., 2021; Kapturowski et al., 2023) as well as Poker (Moravčík et al., 2017; Brown & Sandholm, 2018; 2019). While these techniques work well for games and simulations, their application to practical real-world problems has proven to be more difficult (Dulac-Arnold et al., 2020). This has changed in recent years, where a number of breakthroughs have been achieved by transferring RL policies trained in simulation to the real world (Degrave et al., 2022; Kaufmann et al., 2023) or by successfully applying RL to problems that were traditionally considered supervised problems (Ouyang et al., 2022; Fawzi et al., 2022; Mankowitz et al., 2023). It has long been known that any supervised learning (SL) problem can be reformulated as an RL problem (Barto & Dietterich, 2004) by defining rewards that match the loss function. This idea has not been used much in practice because the advantage of RL has been unclear, and RL problems have been considered to be harder to solve. The key advantage of reinforcement learning over supervised learning is that the optimization objective does not need to be differentiable. To see why this property is important, consider the task of text prediction, at which models like ChatGPT had a lot of success recently. The large language models used in this task are pre-trained using self-supervision (Brown et al., 2020) on a large corpus of internet text, which allows them to generate realistic and linguistically flawless responses to text prompts. However, self-supervised models like GPT-3 cannot directly be deployed in products because they are not optimized to predict helpful, honest, and harmless answers (Bai et al., 2022). So far, the most successful technique to address this problem is called reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; Bai et al., 2022; Ouyang et al., 2022) in which human annotators rank outputs of the model and the task is to maximize this ranking. The mapping between the models outputs and a human ranking is not differentiable, hence supervised learning cannot optimize this objective, whereas reinforcement learning techniques can. Recently, RL was also able to claim success in code generation (Mankowitz et al., 2023) by maximizing execution speed of predicted code, discovering new optimization techniques. Execution speed of code can easily be measured, but not computed in a differentiable way.
@@
This recent success of reinforcement learning on real world problems makes it likely that RL techniques will become relevant for the broader machine learning audience. However, the field of RL currently has a large entry barrier, requiring a significant time investment to get started. Seminal work in the field (Schulman et al., 2015; 2016; Bellemare et al., 2017; Haarnoja et al., 2018a) often focuses on rigorous theoretical exposition and typically assumes that the reader is familiar with prior work. Existing textbooks (Sutton & Barto, 2018; François-Lavet et al., 2018) make little assumptions but are extensive in length. Our aim is to provide readers that are familiar with supervised machine learning an easy entry into the field of deep reinforcement learning to facilitate widespread adoption of these techniques. Towards this goal, we skip the typically rather lengthy introduction via tables and Markov decision processes. Instead, we introduce deep reinforcement learning through the intuitive lens of optimization. In only 24 pages, we introduce the reader to all relevant concepts to understand successful modern Deep RL algorithms like proximal policy optimization (PPO) (Schulman et al., 2017) or soft actor-critic (SAC) (Haarnoja et al., 2018a).
@@ 3.1 Value Learning
The most popular idea to maximize rewards without a gradient from the action $a$ to the reward $r$ is value learning (Sutton & Barto, 2018). The key idea of value learning is to directly predict the reward $r$ rather than the action $a$ and to define the policy $\pi$ implicitly. Optimization is carried out by minimizing the difference between the predicted reward and the observed reward using a regression loss.
@@ 3.1.1 Discrete Action Spaces: Q-Learning
Figure 2: Q-Functions. We illustrate the predicted reward of a Q-function for a fixed state. (1(a)) Discrete action space with 5 classes. The best action can be selected by computing all 5 Q-values (sequentially or in parallel). (1(b)) 1-dimensional continuous action space. The maximum value cannot easily be found since the Q-function can only be evaluated at a finite amount of points. Instead, a policy network predicts the action with the highest reward. The policy is improved by following the gradient of the Q-function uphill.
@@ 3.1.2 Continuous Action Spaces: Deterministic Policy Gradients
Note that the $\operatorname*{argmax}$ in Eq. (3) is intractable in settings with continuous actions. A common solution to this problem is to define the policy explicitly as a neural network that predicts the $\operatorname*{argmax}$ function. In this case, the policy $\pi$ is optimized to maximize the output of the Q-function with the following loss:
@@
Figure 3: Optimization of Non-Differentiable Objectives. We compare Q-learning (continuous setting/actor-critic) in (2(a)) to stochastic policy gradients in (2(b)). Note that both Q-learning and stochastic policy gradients do not require differentiation of the environment.
@@
3.1.3 Example: Image Classification with Q-Learning
Q-learning as we have discussed it so far is already sufficient, in principle, to handle most typical machine learning tasks. We will demonstrate this here with a simple image classification problem and the non-differentiable accuracy metric that is frequently used to measure the performance of classification models but that cannot directly be optimized using gradient-based optimization techniques. Using the tools of RL in combination with a static dataset is called Offline RL. Our goal is to train a ResNet50 (He et al., 2016) for image classification on CIFAR-10 (Krizhevsky et al., 2009), optimizing accuracy directly with Q-learning. In this setting, the state is a 32x32 pixel image, the actions are the 10 class labels and the reward objective is the per class accuracy. As the action space is discrete, we use a ResNet50 (He et al., 2016) with 10 output nodes as our Q-function. The accuracy reward is 1 for the correct class and 0 otherwise. The reward labels can therefore be naturally represented as one-hot vectors, just like in supervised learning. Thus, changing the loss function from cross-entropy to mean squared error is the only change necessary to switch from supervised learning to Q-learning. We use the standard hyperparameters of the image classification library timm (version 0.9.7) (Wightman, 2019), except for a 10 times larger learning rate when training with the MSE loss. To classify an image, we select the class for which the Q-ResNet predicts the highest accuracy. We compare the MSE loss to the standard cross entropy (CE) loss without label smoothing on the CIFAR-10 dataset: Loss Validation Accuracy $\uparrow$ Cross-Entropy 95.1 Mean Squared Error 95.4 As evident from the results above, Q-learning achieves similar accuracy to the cross-entropy (CE) loss. While it is well known that classification models can be trained with MSE (Hastie et al., 2009), here we show that this can be viewed as offline Q-learning and leads to policies that maximize accuracy. This may be surprising since CE is used as the de-facto standard loss function for classification, motivated by information theory, while it is considered a surrogate loss for accuracy (Song et al., 2016; Grabocka et al., 2019; Huang et al., 2021). However, as we will see in Section 3.2, the CE loss can be interpreted as another reinforcement learning technique (stochastic policy gradients).
@@ 3.2.1 Example: Image Classification with Stochastic Policy Gradients
Consider again the example of classification, but this time with policy gradients. The reward is accuracy $acc(s,a)$. The actions are classes and the accuracy is one, if $a$ matches the label $a^{\star}$ and zero otherwise: $\nabla_{\pi}J(\pi)=\frac{1}{|\mathcal{S}|}\sum_{(s,a)\in\mathcal{S}}acc(s,a)\nabla_{\pi}\pi(a|s)$ (8) Since all terms with zero accuracy cancel, we can simplify this equation as follows: $\nabla_{\pi}J(\pi)=\frac{1}{|\mathcal{S}|}\sum_{(s,a^{\star})\in\mathcal{S}}\nabla_{\pi}\pi(a^{\star}|s)$ (9) Instead of maximizing $\pi(a^{\star}|s)$, one may choose to minimize the negative $\log$ probability $-\log\pi(a^{\star}|s)$, since the logarithm is a monotonic function and hence does not change the location of the global optimum. Doing so recovers the familiar cross-entropy loss which corresponds to the negative log-likelihood: $L_{\pi}\,=\,\underbrace{-\sum_{(s,a^{\star})\in\mathcal{S}}\log\pi(a^{\star}|s)}_{\text{Cross Entropy}}\,=\,\underbrace{-\log\prod_{(s,a^{\star})\in\mathcal{S}}\pi(a^{\star}|s)}_{\text{Negative Log-Likelihood}}$ (10) We see that accuracy maximization is another motivation for cross-entropy, besides the standard information theoretic derivation (Shore & Johnson, 1980) and maximum likelihood estimation (Bishop, 2006).
@@ 4.1 The Compounding Error Problem
A typical assumption that neural networks require for generalization is that the data they are trained with covers the underlying distribution well. The dataset is assumed to contain independent and identically distributed (i.i.d.) samples from the data generating distribution. Validation sets used to evaluate a model often satisfy this assumption because they are randomly sampled subsets of the whole dataset. In the i.i.d. setting, neural networks typically work well. However, neural networks are known to struggle with out-of-distribution (OOD) data which is very different from the training data (Beery et al., 2018).
In sequential decision making problems, it is very hard to construct datasets whose samples are i.i.d. because of the feedback loop. During inference, the next state a neural network observes depends on the actions it has predicted in past states. Since the network might take different actions than an annotator that collected the dataset, it may visit states very different from those present in the dataset, i.e. the network will encounter out-of-distribution data at test time. Consider for example the problem of autonomous driving. A human annotator may collect a dataset by driving along various routes. As the data collectors are expert drivers, they will always drive near the center of lanes. A neural network trained with such data may make mistakes at test time, deviating from the center of the lane. Its new inputs will now be out-of-distribution, because the human expert drivers did not make such a mistake, typically leading to even worse predictions, compounding errors, until the policy catastrophically fails. In other words, the training data is missing examples of how to recover from mistakes that an annotator does not make.
Figure 4: Compounding Error Problem. Small mistakes lead to OOD states that increase the error.
@@ 4.2 Exploration and Exploitation
By picking the best currently known action, the agent is following the current policy during data collection. This behavior is called exploitation. However, to discover new states, other actions must sometimes be chosen. This is called exploration. In settings with discrete action spaces, exploration can be performed by picking a random action with a certain probability $\epsilon$ and the best known action otherwise. This strategy is called $\epsilon$-greedy. The $\epsilon$ parameter regulates the exploration-exploitation tradeoff. Additionally, the $\epsilon$ hyperparameter can be decayed towards zero over the course of training, analogously to learning rate schedules in supervised learning. In settings with continuous action spaces, noise can be added to the predicted action instead. Typically, Gaussian noise with mean zero and a standard deviation that is a hyperparameter analogously to the $\epsilon$ value (Eberhard et al., 2023) is used.
@@
In environments where random actions induce catastrophic failure, performing random actions may prevent the agent from reaching states far away from the start, rendering data collection with $\epsilon$-greedy brittle. One approach to mitigate this problem is to learn the amount of noise to be applied. With policies that output a probability distribution, the action can be sampled from this distribution. Algorithms using this idea typically add an entropy bonus to their training objective, so that the network is encouraged to explore but only if exploration doesn’t impact the expected performance too much (Schulman et al., 2017; Haarnoja et al., 2018a; b). For tasks with deterministic outputs, a learnable amount of Gaussian noise can be added to the hidden units in the linear layers instead (Fortunato et al., 2018; Hessel et al., 2018).
@@
The exploration-exploitation tradeoff may be somewhat of a misleading name in the context of RL. Unlike animals, that need to learn at test time, RL agents typically have a training phase where the agent is trying to learn and performance is not relevant. At test time, the trained agent does not learn anymore and only exploits the best known option. Since performance does not matter during training, it may seem beneficial to only explore and never exploit. This, however, is not a good strategy because the agent is unlikely to reach interesting states far away from the start while only performing random actions. The purpose of exploitation during training is to return to previously visited promising states, such that the network can explore from relevant states different from the initial state. Hence, it might be better to speak of the exploration-return tradeoff in the context of RL. Instead of using exploitation to return, one may also choose to simply reset a simulator to a particular state. Ecoffet et al. (2021) has used this idea by keeping a buffer of all visited states. They collected data by simply resetting the simulator to previously visited states and performing random exploration. This approach achieved above human level results on all Atari games (Bellemare et al., 2013), which is an important RL benchmark. However, keeping a buffer of all visited states requires strong compression and might not scale to more complex environments.
@@
A particularly difficult challenge for data collection is posed by environments where random actions will not lead to states with different rewards, i.e., because rewards are sparse. For example, consider an environment where the agent is playing chess against a grand-master level chess bot and the reward is +1 for a win, -1 for a loss and 0 otherwise. A policy collecting data by performing random actions will always lose, and hence all data will have the same reward of -1. Reward optimization is prevented because no data with high rewards will be collected, all actions seem equally bad to the optimizer. Various solutions to the problem of sparse rewards have been proposed. One integral technique in chess and other board games is called self-play (Samuel, 1959; Tesauro, 1995; Silver et al., 2017). Instead of playing against a grand-master level opponent from the start, the policy is playing against itself. This has the effect that the difficulty of the task is automatically adjusted to the skill of the current policy. Self-play ensures that both victory and losses are observed, and winning automatically becomes harder as the policy improves. Another option is to incorporate prior knowledge, for example in the form of human expert demonstrations (Silver et al., 2016; Vecerík et al., 2017; Chekroun et al., 2021). Lastly, some approaches maximize different rewards that make optimization easier and guide data collection. Examples include shaped rewards (Ng et al., 1999) which are reward functions that are engineered for that purpose or intrinsic motivation (Aubret et al., 2019) which are objectives that guide data collection towards states that novel or “surprising”. While there are some solutions for particular environments, optimizing environments with sparse rewards is still a subject of ongoing research (Chen et al., 2022).

### v2

@@ 1 Introduction
The field of reinforcement learning (RL) is traditionally viewed as the art of learning by trial and error (Sutton & Barto, 2018). Reinforcement learning methods were historically developed to solve sequential decision making tasks. The core idea is to deploy an untrained model in an environment. This model is called the policy and maps inputs to actions. The policy is then improved by randomly attempting different actions and observing an associated feedback signal, called the reward. Reinforcement learning techniques have demonstrated remarkable success when applied to popular games. For example, RL produced world-class policies in the games of Go (Silver et al., 2016; 2018; Schrittwieser et al., 2020), Chess (Silver et al., 2018; Schrittwieser et al., 2020), Shogi (Silver et al., 2018; Schrittwieser et al., 2020), Starcraft (Vinyals et al., 2019), Stratego (Perolat et al., 2022) and achieved above human level policies in all Atari games (Badia et al., 2020; Ecoffet et al., 2021; Kapturowski et al., 2023) as well as Poker (Moravčík et al., 2017; Brown & Sandholm, 2018; 2019). While these techniques work well for games and simulations, their application to practical real-world problems has proven to be more difficult (Dulac-Arnold et al., 2020). This has changed in recent years, where a number of breakthroughs have been achieved by transferring RL policies trained in simulation to the real world (Bellemare et al., 2020; Degrave et al., 2022; Kaufmann et al., 2023) or by successfully applying RL to problems that were traditionally considered supervised problems (Ouyang et al., 2022; Fawzi et al., 2022; Mankowitz et al., 2023). It has long been known that any supervised learning (SL) problem can be reformulated as an RL problem (Barto & Dietterich, 2004; Jiang et al., 2023) by defining rewards that match the loss function. This idea has not been used much in practice because the advantage of RL has been unclear, and RL problems have been considered to be harder to solve. A key advantage of reinforcement learning over supervised learning is that the optimization objective does not need to be differentiable. To see why this property is important, consider the task of text prediction, at which models like ChatGPT had a lot of success recently. The large language models used in this task are pre-trained using self-supervision (Brown et al., 2020) on a large corpus of internet text, which allows them to generate realistic and linguistically flawless responses to text prompts. However, self-supervised models like GPT-3 cannot directly be deployed in products because they are not optimized to predict helpful, honest, and harmless answers (Bai et al., 2022). So far, the most successful technique to address this problem is called reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; Bai et al., 2022; Ouyang et al., 2022) in which human annotators rank outputs of the model and the task is to maximize this ranking. The mapping between the models outputs and a human ranking is not differentiable, hence supervised learning cannot optimize this objective, whereas reinforcement learning techniques can. Recently, RL was also able to claim success in code generation (Mankowitz et al., 2023) by maximizing execution speed of predicted code, discovering new optimization techniques. Execution speed of code can easily be measured, but not computed in a differentiable way. Derivative-free optimization methods (Hansen, 2016; Frazier, 2018) can also optimize non-differentiable objectives but typically do not scale well to deep neural networks. A second advantage RL has over supervised learning is that algorithms can collect their own data which allows them to discover novel solutions (Silver et al., 2016; Mankowitz et al., 2023), that a static human annotated dataset might not contain.
@@
The recent success of reinforcement learning on real world problems makes it likely that RL techniques will become relevant for the broader machine learning audience. However, the field of RL currently has a large entry barrier, requiring a significant time investment to get started. Seminal work in the field (Schulman et al., 2015; 2016; Bellemare et al., 2017; Haarnoja et al., 2018a) often focuses on rigorous theoretical exposition and typically assumes that the reader is familiar with prior work. Existing textbooks (Sutton & Barto, 2018; François-Lavet et al., 2018) make little assumptions but are extensive in length. Our aim is to provide readers that are familiar with supervised machine learning an easy entry into the field of deep reinforcement learning to facilitate widespread adoption of these techniques. Towards this goal, we skip the typically rather lengthy introduction via tables and Markov decision processes. Instead, we introduce deep reinforcement learning through the intuitive lens of optimization. In only 25 pages, we introduce the reader to all relevant concepts to understand successful modern Deep RL algorithms like proximal policy optimization (PPO) (Schulman et al., 2017) or soft actor-critic (SAC) (Haarnoja et al., 2018a).
@@ 3.1 Value Learning
The most popular idea to maximize rewards without a gradient from the action $a$ to the reward $r$ is value learning (Sutton & Barto, 2018). The key idea of value learning is to directly predict the (expected) return rather than the action $a$ and to define the policy $\pi$ implicitly. Value learning simplifies to predicting the current reward $r$ in environments of length one, as there is no future reward. We generalize value learning to multi-step problems in Section 5. Optimization is carried out by minimizing the difference between the predicted reward and the observed reward using a regression loss.
@@ 3.1.1 Discrete Action Spaces: Q-Learning
Figure 2: Q-Functions. We illustrate the predicted reward of a Q-function for a fixed state. (2(a)) Discrete action space with 5 classes. The best action can be selected by computing all 5 Q-values (sequentially or in parallel). (2(b)) 1-dimensional continuous action space. The maximum value cannot easily be found since the Q-function can only be evaluated at a finite amount of points. Instead, a policy network predicts the action with the highest reward. The policy is improved by following the gradient of the Q-function uphill.
3.1.2 Example: Image Classification with Q-Learning
To demonstrate how Q-learning can be used to optimize a non-differentiable objective, we will use a simple image classification problem and the non-differentiable accuracy metric that is frequently used to measure the performance of classification models but that cannot directly be optimized using gradient-based optimization techniques. Using the optimization techniques of RL in combination with a static dataset is called Offline RL.
Our goal is to train a ResNet50 (He et al., 2016) for image classification on CIFAR-10 (Krizhevsky et al., 2009), optimizing accuracy directly with Q-learning. In this setting, the state is a 32x32 pixel image, the actions are the 10 class labels and the reward objective is the per class accuracy. As the action space is discrete, we use a ResNet50 (He et al., 2016) with 10 output nodes as our Q-function. The accuracy reward is 1 for the correct class and 0 otherwise. The reward labels can therefore be naturally represented as one-hot vectors, just like in supervised learning. Thus, changing the loss function from cross-entropy to mean squared error is the only change necessary to switch from supervised learning to Q-learning. We use the standard hyperparameters of the image classification library timm (version 0.9.7) (Wightman, 2019), except for a 10 times larger learning rate when training with the MSE loss. To classify an image, we select the class for which the Q-ResNet predicts the highest accuracy. We compare the MSE loss to the standard cross entropy (CE) loss without label smoothing on the CIFAR-10 dataset:
Loss Validation Accuracy $\uparrow$
Cross-Entropy 95.1
Mean Squared Error 95.4
As evident from the results above, Q-learning achieves similar accuracy to the cross-entropy (CE) loss. While it is well known that classification models can be trained with MSE (Hastie et al., 2009), here we show that this can be viewed as offline Q-learning and leads to policies that maximize accuracy. This result doesn’t imply that Q-learning is competitive in every classification task. Q-learning can struggle to scale naively to high dimensional action spaces (de Wiele et al., 2020), which is important for classification benchmarks like ImageNet (Deng et al., 2009). It may be surpising that CE achieves similar accuracy on CIFAR-10, since it is motivated by information theory and considered a surrogate loss for accuracy (Song et al., 2016; Grabocka et al., 2019; Huang et al., 2021). However, as we will see in Section 3.2, the CE loss can be interpreted as another reinforcement learning technique (stochastic policy gradients).
@@ 3.1.3 Continuous Action Spaces: Deterministic Policy Gradients
When using Q-learning in settings with continuous action spaces, the problem arises that the $\operatorname*{argmax}$ in Eq. (3) is intractable. A common solution to this problem is to define the policy explicitly as a neural network that predicts the $\operatorname*{argmax}$ function. In this case, the policy $\pi$ is optimized to maximize the output of the Q-function with the following loss:
@@
Figure 3: Optimization of Non-Differentiable Objectives. We compare Q-learning (continuous setting/actor-critic) in (3(a)) to stochastic policy gradients in (3(b)). Note that both Q-learning and stochastic policy gradients do not require differentiation of the environment.
@@ 3.2.1 Example: Image Classification with Stochastic Policy Gradients
Consider again the example of classification, but this time with policy gradients. The reward is accuracy $acc(s,a)$. The actions are classes and the accuracy is one, if $a$ matches the label $a^{\star}$ and zero otherwise:
$\nabla_{\pi}J(\pi)=\frac{1}{|\mathcal{S}|}\sum_{(s,a)\in\mathcal{S}}acc(s,a)\nabla_{\pi}\pi(a|s)$ (8)
Since all terms with zero accuracy cancel, we can simplify this equation as follows:
$\nabla_{\pi}J(\pi)=\frac{1}{|\mathcal{S}|}\sum_{(s,a^{\star})\in\mathcal{S}}\nabla_{\pi}\pi(a^{\star}|s)$ (9)
Instead of maximizing $\pi(a^{\star}|s)$, one may choose to minimize the negative $\log$ probability $-\log\pi(a^{\star}|s)$, since the logarithm is a monotonic function and hence does not change the location of the global optimum. Doing so recovers the familiar cross-entropy loss which corresponds to the negative log-likelihood:
$L_{\pi}\,=\,\underbrace{-\sum_{(s,a^{\star})\in\mathcal{S}}\log\pi(a^{\star}|s)}_{\text{Cross Entropy}}\,=\,\underbrace{-\log\prod_{(s,a^{\star})\in\mathcal{S}}\pi(a^{\star}|s)}_{\text{Negative Log-Likelihood}}$ (10)
We see that accuracy maximization is another motivation for cross-entropy, besides the standard information theoretic derivation (Shore & Johnson, 1980) and maximum likelihood estimation (Bishop, 2006).
@@ 4.1 The Compounding Error Problem
A typical assumption that neural networks require for generalization is that the data they are trained with covers the underlying distribution well. The dataset is assumed to contain independent and identically distributed (IID) samples from the data generating distribution. Validation sets used to evaluate a model often satisfy this assumption because they are randomly sampled subsets of the whole dataset. In the IID setting, neural networks typically work well. However, neural networks are known to struggle with Non-IID data which is very different from the training data (Beery et al., 2018; Schölkopf, 2022).
In sequential decision making problems, it is very hard to construct datasets whose samples are IID because of the feedback loop. During inference, the next state a neural network observes depends on the actions it has predicted in past states. Since the network might take different actions than an annotator that collected the dataset, it may visit states very different from those present in the dataset, i.e. the network will encounter out-of-distribution data at test time. Consider for example the problem of autonomous driving. A human annotator may collect a dataset by driving along various routes. As the data collectors are expert drivers, they will always drive near the center of lanes. A neural network trained with such data may make mistakes at test time, deviating from the center of the lane. Its new inputs will now be out-of-distribution, because the human expert drivers did not make such a mistake, typically leading to even worse predictions, compounding errors, until the policy catastrophically fails. In other words, the training data is missing examples of how to recover from mistakes that an annotator does not make.
Figure 4: Compounding Error Problem. Small mistakes lead to Non-IID states that increase the error.
@@ 4.2 Exploration and Exploitation
By picking the best currently known action, the agent is following the current policy during data collection. This behavior is called exploitation. However, to discover new states, other actions must sometimes be chosen. This is called exploration. In settings with discrete action spaces, exploration can be performed by picking a random action with a certain probability $\epsilon$ and the best known action otherwise. This strategy is called $\epsilon$-greedy. The $\epsilon$ parameter regulates the exploration-exploitation tradeoff. Additionally, the $\epsilon$ hyperparameter can be decayed towards zero over the course of training. In settings with continuous action spaces, noise can be added to the predicted action instead. Typically, Gaussian noise with mean zero and a standard deviation that is a hyperparameter analogously to the $\epsilon$ value (Eberhard et al., 2023) is used.
@@
In environments where random actions induce catastrophic failure, performing random actions may prevent the agent from reaching states far away from the start, rendering data collection with $\epsilon$-greedy brittle. One approach to mitigate this problem is to learn the amount of noise to be applied. With policies that output a probability distribution, the action can be sampled from this distribution. Algorithms using this idea typically add an objective to increase the entropy of the policy to their training objective, so that the network is encouraged to explore but only if exploration doesn’t impact the expected performance too much (Schulman et al., 2017; Haarnoja et al., 2018a; b). For tasks with deterministic outputs, a learnable amount of Gaussian noise can be added to the hidden units in the linear layers instead (Fortunato et al., 2018; Hessel et al., 2018).
@@
The exploration-exploitation tradeoff may be somewhat of a misleading name in the context of RL. Unlike animals, that need to learn at test time, RL agents typically have a training phase where the agent is trying to learn and only the final performance matters. At test time, the trained agent does not learn anymore and only exploits the best known option. Since the rewards that the agent obtains during training typically do not matter to the designer of the algorithm, it may seem beneficial to only explore and never exploit such that the policy can learn as much as possible. This, however, is not a good strategy because the agent is unlikely to reach interesting states far away from the start while only performing random actions. The purpose of exploitation during training is to return to previously visited promising states, such that the network can explore from relevant states different from the initial state. Hence, it might be better to speak of the exploration-return tradeoff in the context of RL. Instead of using exploitation to return, one may also choose to simply reset a simulator to a particular state. Ecoffet et al. (2021) has used this idea by keeping a buffer of all visited states. They collected data by simply resetting the simulator to previously visited states and performing random exploration. This approach achieved above human level results on all Atari games (Bellemare et al., 2013), which is an important RL benchmark. However, keeping a buffer of all visited states requires strong compression and might not scale to more complex environments.
@@
A particularly difficult challenge for data collection is posed by environments where random actions will not lead to states with different rewards, i.e., because rewards are sparse. For example, consider an environment where the agent is playing chess against a grand-master level chess bot and the reward is +1 for a win, -1 for a loss and 0 otherwise. A policy collecting data by performing random actions is extremely unlikely to win, and hence all collected data will have the same reward of -1. Reward optimization is prevented because no data with high rewards will be collected, all actions seem equally bad to the optimizer. This is also the case in this example if the agent samples an improved action, since the optimizer will not know that the action was better unless it ultimately led to victory. Various solutions to the problem of sparse rewards have been proposed. One integral technique in chess and other board games is called self-play (Samuel, 1959; Tesauro, 1995; Silver et al., 2017). Instead of playing against a grand-master level opponent from the start, the policy is playing against itself or a past version of itself. This has the effect that the difficulty of the task is automatically adjusted to the skill of the current policy. Self-play ensures that both victory and losses are observed, and winning automatically becomes harder as the policy improves. Another option is to incorporate prior knowledge, for example in the form of human expert demonstrations (Silver et al., 2016; Vecerík et al., 2017; Chekroun et al., 2021). Lastly, some approaches maximize different rewards that make optimization easier and guide data collection. Examples include shaped rewards (Ng et al., 1999) which are reward functions that are engineered for that purpose or intrinsic motivation (Aubret et al., 2019) which are objectives that guide data collection towards states that novel or “surprising”. While there are some solutions for particular environments, optimizing environments with sparse rewards is still a subject of ongoing research (Chen et al., 2022).
Another idea to make sparse reward settings or more generally hard tasks easier to learn is curriculum learning (Selfridge et al., 1985; Krueger & Dayan, 2009; Bengio et al., 2009). In curriculum learning, the environment or data is changed over time, typically increasing the difficulty of the task. The goal is that by learning the easier task first, the policy develops abilities that help it solve the harder task that appear later during training. Curriculum learning can aid data collection when the agent, by solving the easier environments, acquires the ability to reach promising states in the harder environments, that it would not have reached if it started from scratch in the harder environment. Additionally, curriculum learning was also empirically observed to improve optimization (Bengio et al., 2009). Designing a curriculum of increasingly harder environments requires domain knowledge and engineering but can be a crucial technique to further improve the performance of state-of-the-art methods. An example would be legged locomotion (Lee et al., 2020; Miki et al., 2022), a task where a legged robot needs to learn how to walk, where the difficulty of the terrain as well as external disturbances applied to the robot are gradually increased, adapting the difficulty of the environment to the ability of the policy. Automating the design of curricula is currently an active area of research (Leibo et al., 2019; Portelas et al., 2020; Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022; Li et al., 2024) called automatic curriculum learning. Self-play, introduced in the last paragraph, can also be viewed as an automatic curriculum learning method, because it can be used without much domain knowledge in environments that feature competitive play.